OpenStudy (anonymous):

5-2x – 3+x A. 2 – x B. 1 C. X – 2 D. 2 – 2x

jimthompson5910 (jim_thompson5910):

-2x + x = ??

OpenStudy (anonymous):

what?

Parth (parthkohli):

I suppose that the question mentions \((5 - 2x) - (3 + x)\).

Parth (parthkohli):

If so, you should first distribute the minus sign.

OpenStudy (anonymous):

I typed the WHOLE question

OpenStudy (anonymous):

this IS the whole problem

Parth (parthkohli):

Is the question what I mentioned?

Parth (parthkohli):

There are two possibilities for your question, Amy.

jimthompson5910 (jim_thompson5910):

I'm just asking what -2x plus x is

OpenStudy (anonymous):

No i typed the whole question

jimthompson5910 (jim_thompson5910):

That's part of the question, i know

Parth (parthkohli):

OK, well, the spaces confused me a little... but still. :)

Parth (parthkohli):

Continue with what Jim says.

OpenStudy (anonymous):

I just need the answer can some one please help me? :D

jimthompson5910 (jim_thompson5910):

Alright, what is -2 + 1 ?

jimthompson5910 (jim_thompson5910):

I'll help you get to the answer

OpenStudy (anonymous):

5-2x – 3+x 2 - 2x + x = 2 - x Which is choice A.

Parth (parthkohli):

Jim takes $1 for an answer. :)

OpenStudy (anonymous):

Thanks Fall12-13